OpenStudy (chosenmatt):

7x = –2y A. B. C. D.

OpenStudy (anonymous):

Fill in the "options" ??

OpenStudy (anonymous):

what is the answers for A ,B,C,D??

OpenStudy (anonymous):

What are the options provided for "A,B,C,D"?

OpenStudy (chosenmatt):

A-14Y B-9Y C-2Y/7 D -7/2Y

OpenStudy (anonymous):

I think you are tying to figure out what is \(x\) in terms of \(y\)..

OpenStudy (anonymous):

\[7x = -2y \] Divide both the sides by \(7\).. and tell me what did you get?

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

-0.28

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

@chosenmatt getting what I asked?

OpenStudy (chosenmatt):

I got 7 then 5?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\(7x = -2y\) Divide by \(7\) both sides: \[\frac{7x}{7} = \frac{-2y}{7} \\ \frac{\cancel{7}x}{\cancel{7}} = \frac{-2y}{7} \\ x = \frac{-2y}{7}\]

OpenStudy (chosenmatt):

OH!! THX XD